# Pairs of knot invariants

### v2

@@ Abstract.
Let $\alpha$ be a map from the set of all knot types ${\mathcal{K}}$ to a set $X$. Let $\beta$ be a map from ${\mathcal{K}}$ to a set $Y$. We define the relation between $\alpha$ and $\beta$ to be the image of a map $(\alpha,\beta)$ from ${\mathcal{K}}$ to $X\times Y$ sending an element $K$ of ${\mathcal{K}}$ to $(\alpha(K),\beta(K))$. We determine the relations between $\alpha$ and $\beta$ for certain $\alpha$ and $\beta$ such as crossing number, unknotting number, bridge number, braid index, genus and canonical genus.
@@
pair of invariants, pair of knot invariants, relation between invariants, relation between knot invariants, crossing number, unknotting number, bridge number, braid index, genus of knots, canonical genus, delta-unknotting number
@@ 1. Introduction
From now on we restrict our attention to knot invariants. We note that the relation between two knot invariants defined above is a natural concept and implicitly used for a long time. However, as far as the author knows, the explicit definition is not given before. Prior to this concept for a pair of knot invariants, a concept for a single knot invariant exists. Let $X$ be a set and $\alpha:{\mathcal{K}}\to X$ a knot invariant. Then the determination of the image $\alpha({\mathcal{K}})\subset X$ is a fundamental problem. In many cases it is easily determined. For example, let $\Delta:{\mathcal{K}}\to{\mathbb{Z}}[t^{\pm 1}]$ be the Alexander polynomial. It is known that
@@
Let $R$ be a subset of ${\mathbb{R}}$. A map $f:{\mathcal{K}}\to R$ is said to be crossing-order $\leq n$ if there exist real numbers $A_{0},A_{1},\cdots,A_{n}$ such that $|f(K)|\leq A_{0}+A_{1}c(K)+\cdots+A_{n}c(K)^{n}$ for every $K\in{\mathcal{K}}$. Then we see that $a_{2}$ is not crossing-order $\leq 1$. It is shown in [35] that ${|a_{2}(K)|\leq\frac{1}{8}c(K)^{2}}$ for every knot $K$. Therefore $a_{2}$ is crossing-order $\leq 2$. It is shown in [19, Proposition 3.6] that ${a_{2}(K)\leq\frac{1}{8}(c(K)^{2}-1)}$ for every non-trivial knot $K$. It is also remarked in [6, 14.3] that a Vassiliev invariant of order $\leq n$ has crossing-order $\leq n$.
@@
As mentioned above $g(K)\leq g_{c}(K)$ for every oriented knot $K$. However the relation between canonical genus and genus is slightly different from that of braid index minus one and bridge number minus one. It is shown in [43, Theorem 1.1] that no knot has $g_{c}=2$ and $g=1$. The following theorem is implicitly stated without proof in [43].
@@
We mainly consider crossing number, unknotting number, braid index minus one, bridge number minus one, genus and canonical genus. There are $15$ unordered pairs of them. We note that $12$ of them are decided as above. The rest are $(u,g)$, $(u,g_{c})$ and $(g_{c},{\rm bridge}-1)$. It seems to us that they are a bit difficult to determine. To the best of our knowledge, no knot $K$ with $g(K)=1$ and $u(K)\geq 4$ is known. It is only known that $g(P(3,3,3))=1$ and $u(P(3,3,3))=3$ [34]. See also [22]. It is shown in [44, Theorem 2.1] that a canonical genus $1$ knot is a $2$-bridge knot with Conway notation $C(2p,2q)$, or a pretzel knot $P(2p+1,2q+1,2r+1)$. Thus there exists no knot $K$ of $g_{c}(K)=1$ and ${\rm bridge}(K)\geq 4$. In general we have the following theorem.
@@ 2. Crossing number and unknotting number, genus or canonical genus
Therefore $u(P(2p+1,-2,2q+1))\leq i$. Let $S(K)$ be the Rasmussen invariant of an oriented knot $K$ defined in [36]. Suppose that $D$ is a positive diagram of $K$. Then it is shown in [36] that $S(K)=c(D)-s(D)+1$. Since the standard diagram of $P(2p+1,-2,2q+1)$ is a positive $3$-braid diagram, we have $S(P(2p+1,-2,2q+1))=(2i+2)-3+1=2i$. Then by [36] we have $u(P(2p+1,-2,2q+1))\geq i$. Thus we have $u(P(2p+1,-2,2q+1))=i$. Therefore we have $c(P(2p+1,-2,2q+1))\geq 2i+1$. Suppose $c(P(2p+1,-2,2q+1))=2i+1$. Then by [45, Theorem 1.4] we see that $P(2p+1,-2,2q+1)$ is a torus knot $T(2,2i+1)$. It is shown in [42] that
@@ 3. Crossing number and braid index minus one
Since $(c,{\rm braid}-1)(0_{1})=(0,0)$, we have $(0,0)\in(c,{\rm braid}-1)({\mathcal{R}}_{0})$. The braid index of a $2$-bridge knot is determined in [28, Theorem B]. We note that negative continued fractions are used in [28, Theorem B]. Then we see
@@ 5. Crossing number and delta-unknotting number
Let ${\mathcal{S}}$ (resp. ${\mathcal{M}}$) be a subset of $\{1,\cdots,m\}$ such that $p_{i}$ is a preimage of a self crossing of $\alpha$ (resp. mutual crossing of $\alpha$ and $\beta$) if and only if $i\in{\mathcal{S}}$ (resp. $i\in{\mathcal{M}}$). We note that $K$ is ambient isotopic to a band sum of a Hopf link and $K^{\prime}$ where the Hopf link and bands are contained in a small neighbourhood of $x_{0}$ in ${\mathbb{S}}^{3}$ as illustrated in Figure 5.1. Then we lift up the Hopf link and bands so that they form an overbridge. Two vertical parts of the bands are denoted by $b_{1}$ and $b_{2}$ as illustrated in Figure 5.2. The band $b_{1}$ is attached to $K^{\prime}$ at the point $q_{0}$ and the band $b_{2}$ is attached to $K^{\prime}$ at the point $r_{m}$. We denote this form of $K$ by $K(q_{0},r_{m})$. For each $i\in\{0,1,\cdots,m\}$, we fix $i$ and consider the following deformation. We slide $b_{1}$, keeping it vertical, along the orientation of $\alpha$ to $q_{i}$. When $b_{1}$ encounter an over-crossing, $b_{1}$ passes through it by an ambient isotopy. When $b_{1}$ encounter an under-crossing, we perform a “clasp leaps over a hurdle move” illustrated in Figure 5.3 so that $b_{1}$ passes through the under-crossing. This move is realized by an application of delta move [25, Claim 1.1][46, Lemma 2.2]. In fact these two moves are equivalent as local move [47, Examples]. The horizontal parts of bands and a Hopf link are deformed up to ambient isotopy following the move of $b_{1}$. We denote the knots corresponding to the deformation so far by $K(q_{0},r_{m}),K(q_{1},r_{m}),\cdots,K(q_{i},r_{m})$. After $b_{1}$ is moved to $q_{i}$, we slide $b_{2}$, keeping it vertical, along the opposite orientation of $\alpha$ to $r_{i}$, up to ambient isotopy and applications of delta move just as above. We denote the knots corresponding to this deformation starting from $K(q_{i},r_{m})$ by $K(q_{i},r_{m-1}),K(q_{i},r_{m-2}),\cdots,K(q_{i},r_{i})$. The total number of applications of delta move so far is $|\alpha\cap\{x_{1}^{-},\cdots,x_{n}^{-}\}|=|{\mathcal{U}}|$. Namely we have $K(q_{i},r_{i})$ from $K=K(q_{0},r_{m})$ by $|{\mathcal{U}}|$-times application of delta move. Since $|{\mathcal{U}}|\leq\displaystyle{\frac{m}{2}\leq\frac{n}{2}=\frac{c(D)-1}{2}}$, we have $K(q_{i},r_{i})$ from $K$ by at most ${\left\lfloor\frac{c(D)-1}{2}\right\rfloor}$-times applications of delta move. In $K(q_{i},r_{i})$, a Hopf link and bands form a twist knot or a trivial knot as a factor knot connected summed to $K^{\prime}$ at $q_{i}$ and $r_{i}$. Namely $K(q_{i},r_{i})$ is a connected sum of a twist knot or a trivial knot and $K^{\prime}$. The knot type of this factor knot is determined by the total number of full-twists on the bands. This number is equal to the linking number of a $2$-component link obtained by smoothing any one of the two crossings of a Hopf link. In general, we denote the linking number of a $2$-component link obtained from the knot $K(q_{s},r_{t})$ by smoothing any one of the two crossings of a Hopf link by $l(s,t)$. In the following we will show
@@
Let $D$ be a knot diagram. We denote the minimal number of crossings of $D$ changing them turns $D$ to a diagram of a trivial knot by $u(D)$. It is well-known that ${u(D)\leq\frac{c(D)-1}{2}}$ if $c(D)>0$. See for example [45][33, Proposition 2.1]. Moreover we have the following lemma.
@@
Proof. Let $x_{0}$ be a crossing of $D$ that is not outermost. Let $C_{1}$ and $C_{2}$ be closed curves obtained from $D$ by smoothing $x_{0}$. Then both $C_{1}$ and $C_{2}$ are not simple closed curves. Let $x_{1}$ and $x_{2}$ be crossings of $C_{1}$ and $C_{2}$ respectively. Let $U$ be the underlying projection of $D$. Then we see that in the chord diagram of $U$, the chords corresponding to $x_{0}$, $x_{1}$ and $x_{2}$ are mutually parallel. Then we see that the trivializing number $t(U)$ of $U$ defined in [12] is less than or equal to $c(D)-3$ [15, Lemma 3.7]. It is implicitly shown in the proof of [13, Proposition 1] that ${u(D)\leq\frac{t(U)}{2}}$. Therefore we have ${u(D)\leq\frac{c(D)-3}{2}}$. This completes the proof. $\Box$
@@
holds. Let $K$ be a knot with $c(K)\geq 5$. First suppose that $K=T(2,p)$ for some odd number $p$ with $|p|\geq 5$. It is known that $u_{\Delta}(T(2,p))=\displaystyle{\frac{p^{2}-1}{8}}$ [30, Corollary 2.2]. Since $c(T(2,p))=|p|$ it is sufficient to show
@@
Since $|p|^{2}-4|p|-5=(|p|+1)(|p|-5)$ we have shown it. Next suppose that $K$ is not $T(2,p)$ for any odd number $p$. Let $D$ a diagram of $K$ with $c(D)=c(K)$. Then by [45, Theorem 1.4 (1)] we have $u(D)\leq\displaystyle{\frac{c(D)-2}{2}}$. Suppose that there is a crossing of $D$ that is not outermost. Then we have ${u(D)\leq\frac{c(D)-3}{2}}$ by Lemma 5-2. Then by Lemma 5-1 we have
@@ 7. Canonical genus and genus
Proof of Theorem 1-12. As stated in Section 1, it is shown in [43, Theorem 1.1] that no knot has $g_{c}=2$ and $g=1$. Since $g(K)\leq g_{c}(K)$ for every oriented knot $K$, $(g_{c},g)(0_{1})=(0,0)$ and $g(K)=0$ if and only if $K$ is a trivial knot, we have
@@
Let $a$ be an integer with $a\geq 3$. Let $K$ be a $2$-bridge knot with $c(K)=a$ and $D(K)$ a Whitehead double of $K$. Then $c_{g}(D(K))=c(K)=a$ [29, Theorem 1]. Let $\delta(J)$ be the maximal degree in $z$ of the HOMFLY-PT polynomial $P_{J}(v,z)$ of a knot $J$. It is shown in [24] that ${g_{c}(J)\geq\frac{\delta(J)}{2}}$. It is also shown in [29] that $\delta(D(K))=2c(K)=2a$. Since $\delta$ is additive under connected sum of knots and $\delta({3_{1}}^{*})=2$, we have, for a non-negative integer $b$,

### v3

@@ Abstract.
Let $\alpha$ be a map from the set of all knot types ${\mathcal{K}}$ to a set $X$. Let $\beta$ be a map from ${\mathcal{K}}$ to a set $Y$. We define the relation between $\alpha$ and $\beta$ to be the image of a map $(\alpha,\beta)$ from ${\mathcal{K}}$ to $X\times Y$ sending an element $K$ of ${\mathcal{K}}$ to $(\alpha(K),\beta(K))$. We determine the relations between $\alpha$ and $\beta$ for certain $\alpha$ and $\beta$ such as crossing number, unknotting number, bridge number, braid index, genus and canonical genus. This is a study of geography problem in knot theory.
@@
pair of invariants, pair of knot invariants, relation between invariants, relation between knot invariants, geography, botany, crossing number, unknotting number, bridge number, braid index, genus of knots, canonical genus, delta-unknotting number
@@ 1. Introduction
The study of $(\alpha,\beta)({\mathcal{T}})$ is called geography problem in some areas of mathematics. It begins with [38] where the pair $(c_{1}^{2},c_{2})$ of Chern numbers of a surface of general type is considered. The study of $(\alpha,\beta)^{-1}(x,y)$ for an element $(x,y)$ of $X\times Y$ is called botany problem. In this paper we study geography problem for certain pairs of knot invariants.
From now on we restrict our attention to knot invariants. We note that the relation between two knot invariants defined above is a natural concept and implicitly used for a long time. Prior to this concept for a pair of knot invariants, a concept for a single knot invariant exists. Let $X$ be a set and $\alpha:{\mathcal{K}}\to X$ a knot invariant. Then the determination of the image $\alpha({\mathcal{K}})\subset X$ is a fundamental problem. In many cases it is easily determined. For example, let $\Delta:{\mathcal{K}}\to{\mathbb{Z}}[t^{\pm 1}]$ be the Alexander polynomial. It is known that
@@
These studies of knots with ${u(K)=\frac{1}{2}(c(K)-1)}$ or ${u(K)=\frac{1}{2}(c(K)-2)}$ are unified botany problems in knot theory. Some botany problems in knot theory are already studied. For example, genus one unknotting number one knots are doubled knots [23] [45], genus one three-bridge knots are pretzel knots [10], genus one tunnel number one knots are determined in [44], and unknotting number one two-bridge knots are determined in [22].
@@
Let $R$ be a subset of ${\mathbb{R}}$. A map $f:{\mathcal{K}}\to R$ is said to be crossing-order $\leq n$ if there exist real numbers $A_{0},A_{1},\cdots,A_{n}$ such that $|f(K)|\leq A_{0}+A_{1}c(K)+\cdots+A_{n}c(K)^{n}$ for every $K\in{\mathcal{K}}$. Then we see that $a_{2}$ is not crossing-order $\leq 1$. It is shown in [39] that ${|a_{2}(K)|\leq\frac{1}{8}c(K)^{2}}$ for every knot $K$. Therefore $a_{2}$ is crossing-order $\leq 2$. It is shown in [20, Proposition 3.6] that ${a_{2}(K)\leq\frac{1}{8}(c(K)^{2}-1)}$ for every non-trivial knot $K$. It is also remarked in [6, 14.3] that a Vassiliev invariant of order $\leq n$ has crossing-order $\leq n$.
@@
As mentioned above $g(K)\leq g_{c}(K)$ for every oriented knot $K$. However the relation between canonical genus and genus is slightly different from that of braid index minus one and bridge number minus one. It is shown in [49, Theorem 1.1] that no knot has $g_{c}=2$ and $g=1$. The following theorem is implicitly stated without proof in [49].
@@
We mainly consider crossing number, unknotting number, braid index minus one, bridge number minus one, genus and canonical genus. There are $15$ unordered pairs of them. We note that $12$ of them are decided as above. The rest are $(u,g)$, $(u,g_{c})$ and $(g_{c},{\rm bridge}-1)$. It seems to us that they are a bit difficult to determine. To the best of our knowledge, no knot $K$ with $g(K)=1$ and $u(K)\geq 4$ is known. It is only known that $g(P(3,3,3))=1$ and $u(P(3,3,3))=3$ [37]. See also [25]. It is shown in [50, Theorem 2.1] that a canonical genus $1$ knot is a $2$-bridge knot with Conway notation $C(2p,2q)$, or a pretzel knot $P(2p+1,2q+1,2r+1)$. Thus there exists no knot $K$ of $g_{c}(K)=1$ and ${\rm bridge}(K)\geq 4$. In general we have the following theorem.
@@ 2. Crossing number and unknotting number, genus or canonical genus
Therefore $u(P(2p+1,-2,2q+1))\leq i$. Let $S(K)$ be the Rasmussen invariant of an oriented knot $K$ defined in [40]. Suppose that $D$ is a positive diagram of $K$. Then it is shown in [40] that $S(K)=c(D)-s(D)+1$. Since the standard diagram of $P(2p+1,-2,2q+1)$ is a positive $3$-braid diagram, we have $S(P(2p+1,-2,2q+1))=(2i+2)-3+1=2i$. Then by [40] we have $u(P(2p+1,-2,2q+1))\geq i$. Thus we have $u(P(2p+1,-2,2q+1))=i$. Therefore we have $c(P(2p+1,-2,2q+1))\geq 2i+1$. Suppose $c(P(2p+1,-2,2q+1))=2i+1$. Then by [51, Theorem 1.4] we see that $P(2p+1,-2,2q+1)$ is a torus knot $T(2,2i+1)$. It is shown in [48] that
@@ 3. Crossing number and braid index minus one
Since $(c,{\rm braid}-1)(0_{1})=(0,0)$, we have $(0,0)\in(c,{\rm braid}-1)({\mathcal{R}}_{0})$. The braid index of a $2$-bridge knot is determined in [31, Theorem B]. We note that negative continued fractions are used in [31, Theorem B]. Then we see
@@ 5. Crossing number and delta-unknotting number
Let ${\mathcal{S}}$ (resp. ${\mathcal{M}}$) be a subset of $\{1,\cdots,m\}$ such that $p_{i}$ is a preimage of a self crossing of $\alpha$ (resp. mutual crossing of $\alpha$ and $\beta$) if and only if $i\in{\mathcal{S}}$ (resp. $i\in{\mathcal{M}}$). We note that $K$ is ambient isotopic to a band sum of a Hopf link and $K^{\prime}$ where the Hopf link and bands are contained in a small neighbourhood of $x_{0}$ in ${\mathbb{S}}^{3}$ as illustrated in Figure 5.1. Then we lift up the Hopf link and bands so that they form an overbridge. Two vertical parts of the bands are denoted by $b_{1}$ and $b_{2}$ as illustrated in Figure 5.2. The band $b_{1}$ is attached to $K^{\prime}$ at the point $q_{0}$ and the band $b_{2}$ is attached to $K^{\prime}$ at the point $r_{m}$. We denote this form of $K$ by $K(q_{0},r_{m})$. For each $i\in\{0,1,\cdots,m\}$, we fix $i$ and consider the following deformation. We slide $b_{1}$, keeping it vertical, along the orientation of $\alpha$ to $q_{i}$. When $b_{1}$ encounter an over-crossing, $b_{1}$ passes through it by an ambient isotopy. When $b_{1}$ encounter an under-crossing, we perform a “clasp leaps over a hurdle move” illustrated in Figure 5.3 so that $b_{1}$ passes through the under-crossing. This move is realized by an application of delta move [28, Claim 1.1][52, Lemma 2.2]. In fact these two moves are equivalent as local move [53, Examples]. The horizontal parts of bands and a Hopf link are deformed up to ambient isotopy following the move of $b_{1}$. We denote the knots corresponding to the deformation so far by $K(q_{0},r_{m}),K(q_{1},r_{m}),\cdots,K(q_{i},r_{m})$. After $b_{1}$ is moved to $q_{i}$, we slide $b_{2}$, keeping it vertical, along the opposite orientation of $\alpha$ to $r_{i}$, up to ambient isotopy and applications of delta move just as above. We denote the knots corresponding to this deformation starting from $K(q_{i},r_{m})$ by $K(q_{i},r_{m-1}),K(q_{i},r_{m-2}),\cdots,K(q_{i},r_{i})$. The total number of applications of delta move so far is $|\alpha\cap\{x_{1}^{-},\cdots,x_{n}^{-}\}|=|{\mathcal{U}}|$. Namely we have $K(q_{i},r_{i})$ from $K=K(q_{0},r_{m})$ by $|{\mathcal{U}}|$-times application of delta move. Since $|{\mathcal{U}}|\leq\displaystyle{\frac{m}{2}\leq\frac{n}{2}=\frac{c(D)-1}{2}}$, we have $K(q_{i},r_{i})$ from $K$ by at most ${\left\lfloor\frac{c(D)-1}{2}\right\rfloor}$-times applications of delta move. In $K(q_{i},r_{i})$, a Hopf link and bands form a twist knot or a trivial knot as a factor knot connected summed to $K^{\prime}$ at $q_{i}$ and $r_{i}$. Namely $K(q_{i},r_{i})$ is a connected sum of a twist knot or a trivial knot and $K^{\prime}$. The knot type of this factor knot is determined by the total number of full-twists on the bands. This number is equal to the linking number of a $2$-component link obtained by smoothing any one of the two crossings of a Hopf link. In general, we denote the linking number of a $2$-component link obtained from the knot $K(q_{s},r_{t})$ by smoothing any one of the two crossings of a Hopf link by $l(s,t)$. In the following we will show
@@
Let $D$ be a knot diagram. We denote the minimal number of crossings of $D$ changing them turns $D$ to a diagram of a trivial knot by $u(D)$. It is well-known that ${u(D)\leq\frac{c(D)-1}{2}}$ if $c(D)>0$. See for example [51][36, Proposition 2.1]. Moreover we have the following lemma.
@@
Proof. Let $x_{0}$ be a crossing of $D$ that is not outermost. Let $C_{1}$ and $C_{2}$ be closed curves obtained from $D$ by smoothing $x_{0}$. Then both $C_{1}$ and $C_{2}$ are not simple closed curves. Let $x_{1}$ and $x_{2}$ be crossings of $C_{1}$ and $C_{2}$ respectively. Let $U$ be the underlying projection of $D$. Then we see that in the chord diagram of $U$, the chords corresponding to $x_{0}$, $x_{1}$ and $x_{2}$ are mutually parallel. Then we see that the trivializing number $t(U)$ of $U$ defined in [13] is less than or equal to $c(D)-3$ [16, Lemma 3.7]. It is implicitly shown in the proof of [14, Proposition 1] that ${u(D)\leq\frac{t(U)}{2}}$. Therefore we have ${u(D)\leq\frac{c(D)-3}{2}}$. This completes the proof. $\Box$
@@
holds. Let $K$ be a knot with $c(K)\geq 5$. First suppose that $K=T(2,p)$ for some odd number $p$ with $|p|\geq 5$. It is known that $u_{\Delta}(T(2,p))=\displaystyle{\frac{p^{2}-1}{8}}$ [33, Corollary 2.2]. Since $c(T(2,p))=|p|$ it is sufficient to show
@@
Since $|p|^{2}-4|p|-5=(|p|+1)(|p|-5)$ we have shown it. Next suppose that $K$ is not $T(2,p)$ for any odd number $p$. Let $D$ a diagram of $K$ with $c(D)=c(K)$. Then by [51, Theorem 1.4 (1)] we have $u(D)\leq\displaystyle{\frac{c(D)-2}{2}}$. Suppose that there is a crossing of $D$ that is not outermost. Then we have ${u(D)\leq\frac{c(D)-3}{2}}$ by Lemma 5-2. Then by Lemma 5-1 we have
@@ 7. Canonical genus and genus
Proof of Theorem 1-12. As stated in Section 1, it is shown in [49, Theorem 1.1] that no knot has $g_{c}=2$ and $g=1$. Since $g(K)\leq g_{c}(K)$ for every oriented knot $K$, $(g_{c},g)(0_{1})=(0,0)$ and $g(K)=0$ if and only if $K$ is a trivial knot, we have
@@
Let $a$ be an integer with $a\geq 3$. Let $K$ be a $2$-bridge knot with $c(K)=a$ and $D(K)$ a Whitehead double of $K$. Then $c_{g}(D(K))=c(K)=a$ [32, Theorem 1]. Let $\delta(J)$ be the maximal degree in $z$ of the HOMFLY-PT polynomial $P_{J}(v,z)$ of a knot $J$. It is shown in [27] that ${g_{c}(J)\geq\frac{\delta(J)}{2}}$. It is also shown in [32] that $\delta(D(K))=2c(K)=2a$. Since $\delta$ is additive under connected sum of knots and $\delta({3_{1}}^{*})=2$, we have, for a non-negative integer $b$,
@@ 9. Future directions
Acknowledgments
The author is grateful to the referee for letting him know geography and botany problems in mathematics.